# Skeleton-Based Human Action Recognition with Noisy Labels

### v1

@@ Abstract
Understanding human actions from body poses is critical for assistive robots sharing space with humans in order to make informed and safe decisions about the next interaction. However, precise temporal localization and annotation of activity sequences is time-consuming and the resulting labels are often noisy. If not effectively addressed, label noise negatively affects the model’s training, resulting in lower recognition quality. Despite its importance, addressing label noise for skeleton-based action recognition has been overlooked so far. In this study, we bridge this gap by implementing a framework that augments well-established skeleton-based human action recognition methods with label-denoising strategies from various research areas to serve as the initial benchmark. Observations reveal that these baselines yield only marginal performance when dealing with sparse skeleton data. Consequently, we introduce a novel methodology, NoiseEraSAR, which integrates global sample selection, co-teaching, and Cross-Modal Mixture-of-Experts (CM-MOE) strategies, aimed at mitigating the adverse impacts of label noise. Our proposed approach demonstrates better performance on the established benchmark, setting new state-of-the-art standards. The source code for this study will be made accessible at https://github.com/xuyizdby/NoiseEraSAR.
@@ IV-E Cross-Modal Mixture of Expert (CM-MoE)
X-Sub X-View X-Sub X-View X-Sub X-View X-Sub
X-View
CTR-GCN [50]
86.8% 90.9% 83.7% 88.2% 81.7% 85.7% 64.8%
64.6%
CTR-GCN [50] + SOP [16]
83.7% 88.3% 82.0% 87.3% 81.0% 85.6% 69.2%
68.2%
CTR-GCN [50] + NPC [17]
86.6% 90.8% 83.2% 87.4% 82.4% 85.5% 68.9% 70.2%
@@
Let ${S}^{joint}_{c}=\mathbf{N}_{joint}\left(\mathbf{j}_{c}\right)$, ${S}^{bone}_{c}=\mathbf{N}_{bone}\left(\mathbf{b}_{c}\right)$ and $\mathbf{m}_{c}=\mathbf{N}_{motion}\left(\mathbf{m}_{c}\right)$ as the SoftMax score of each expert network, where $\mathbf{j}_{c}$, $\mathbf{b}_{c}$, and $\mathbf{m}_{c}$ are derived from one sample in $D_{c}$. The final prediction result is calculated as follows:
@@ V-B Results
In this section, we compare the performance of our method with two typical existing methods (NPC [17] and SOP [16]), which focus on the noisy label problem in the field of image classification. Table I showcases our method’s performance, subjected to symmetric noise levels ranging from $20\%$ to $80\%$ on the NTU-60 [49] dataset while using CTR-GCN [18] as the backbone.
@@ V-C Qualitative Results
To assess our method against two other approaches, we analyzed four distinct class samples: Drinking water, Typing on a keyboard, Shaking hands, and Shaking head, as shown in Figure 3. Both SOP and NPC methods display less discriminative power, often ranking the correct class within the top $5$ but confusing it with similar classes—for example, NPC mixes up phone call with drinking water.
@@ VI Conclusion
In this paper, we open the vistas for skeleton-based human action recognition under noisy label scenarios. We have constructed a benchmark by using three well-established label-denoising approaches leveraged in other research domains and implemented them with GCN backbones. We propose a new method, NoiseEraSAR, by using co-teaching for multi-modal streams, global sample selection, and CM-MOE. Our proposed approach achieves promising results on the NTU-60 dataset under different label noisy ratios and different evaluation manners. This work will help to achieve a more robust skeleton-based human action recognition model that can serve robot-based human daily life assistance when training with data containing low-quality labels and will contribute to the community.

### v2

@@ Abstract
Understanding human actions from body poses is critical for assistive robots sharing space with humans in order to make informed and safe decisions about the next interaction. However, precise temporal localization and annotation of activity sequences is time-consuming and the resulting labels are often noisy. If not effectively addressed, label noise negatively affects the model’s training, resulting in lower recognition quality. Despite its importance, addressing label noise for skeleton-based action recognition has been overlooked so far. In this study, we bridge this gap by implementing a framework that augments well-established skeleton-based human action recognition methods with label-denoising strategies from various research areas to serve as the initial benchmark. Observations reveal that these baselines yield only marginal performance when dealing with sparse skeleton data. Consequently, we introduce a novel methodology, NoiseEraSAR, which integrates global sample selection, co-teaching, and Cross-Modal Mixture-of-Experts (CM-MOE) strategies, aimed at mitigating the adverse impacts of label noise. Our proposed approach demonstrates better performance on the established benchmark, setting new state-of-the-art standards. The source code for this study is accessible at https://github.com/xuyizdby/NoiseEraSAR.
@@ IV-E Cross-Modal Mixture of Expert (CM-MoE)
X-Sub X-View X-Sub X-View X-Sub X-View X-Sub X-View
CTR-GCN [50] 86.8% 90.9% 83.7% 88.2% 81.7% 85.7% 64.8% 64.6%
CTR-GCN [50] + SOP [16] 83.7% 88.3% 82.0% 87.3% 81.0% 85.6% 69.2% 68.2%
CTR-GCN [50] + NPC [17] 86.6% 90.8% 83.2% 87.4% 82.4% 85.5% 68.9% 70.2%
@@
Let ${S}^{joint}_{c}=\mathbf{N}_{joint}\left(\mathbf{j}_{c}\right)$, ${S}^{bone}_{c}=\mathbf{N}_{bone}\left(\mathbf{b}_{c}\right)$ and ${S}^{motion}_{c}=\mathbf{N}_{motion}\left(\mathbf{m}_{c}\right)$ as the SoftMax score of each expert network, where $\mathbf{j}_{c}$, $\mathbf{b}_{c}$, and $\mathbf{m}_{c}$ are derived from one sample in $D_{c}$. The final prediction result is calculated as follows:
@@ V-B Results
In this section, we compare the performance of our method with two typical existing methods (NPC [17] and SOP [16]), which focus on the noisy label problem in the field of image classification. Table I showcases our method’s performance, subjected to symmetric noise levels ranging from $20\%$ to $80\%$ on the NTU-60 [49] dataset while using CTR-GCN [18] as the backbone. We randomly choose $4$ different noise ratios among $0.0\%$ to $100.0\%$ to set up this benchmark, encompassing easy, middle, and hard levels.
@@ V-C Qualitative Results
To further assess our method against two other approaches, we select samples from four action classes that are easily confused by other categories when recognized by the human eye: Drinking water, Typing on a keyboard, Shaking hands, and Shaking head, as shown in Figure 3. Both SOP and NPC methods display less discriminative power, often ranking the correct class within the top $5$ but confusing it with similar classes—for example, NPC mixes up phone call with drinking water.
@@ VI Conclusion
In this paper, we explore skeleton-based human action recognition with noisy labels. We created a benchmark using three label-denoising methods applied to GCN backbones. Our new method, NoiseEraSAR, employs co-teaching for multi-modal streams, global sample selection, and CM-MOE. NoiseEraSAR shows promising results on the NTU-60 dataset across various noisy label ratios and evaluation methods. This work aims to enhance robust skeleton-based action recognition models for robot-assisted human activities, even with low-quality training data, benefiting the community.